OpenStudy (maddy1251):

1. Use your equations from Lesson 1 to create a graph of two polynomial functions so that you can compare volumes for different values of x. Compare the volumes. (I can provide these to who needs them) and 2. Use the graph from Question 1 to identify where the volumes are equal. Confirm this value by writing and solving an equation that describes the difference between the volumes of the candle types. Just looking for guidance to help, not someone to do my homework.

OpenStudy (maddy1251):

I am lost and don't know where to begin

OpenStudy (anonymous):

we are all lost

OpenStudy (anonymous):

since we don't have question one

OpenStudy (clalgee):

@Maddy1251 Post the graph :)

OpenStudy (maddy1251):

@Clalgee I do not have graphs yet however I have equations.

OpenStudy (luigi0210):

@ganeshie8

OpenStudy (luigi0210):

@Maddy1251 Put in the rest of the information.. gane is the best person for help :P

ganeshie8 (ganeshie8):

Volumes are equal at the point where the two curves intersect ``` (I can provide these to who needs them) ``` where are the polynomials and graphs ?

ganeshie8 (ganeshie8):

for graphing, you may click below link and enter the equations of polynomials https://www.desmos.com/calculator

OpenStudy (maddy1251):

Brick Polynomial equation is V=x^³-x^² and Egyptian polynomial equation is V= 1/3x^2+7x (from what I found)

OpenStudy (maddy1251):

They want us to make a graph, so I do not have one yet. Thats what I am trying to create.

ganeshie8 (ganeshie8):

got you :) for part1, enter below equations in desmos site : y = x^3 - x^2 y = 1/3x^2 + 7x

ganeshie8 (ganeshie8):

wait, the second volume equations doesn't look correct as its degree is just 2 can you double check the equations ?